# Attend or Perish: Benchmarking Attention in Algorithmic Reasoning

### v1

@@ Abstract
Can transformers learn to perform algorithmic tasks reliably across previously unseen input/output domains? While pre-trained language models show solid accuracy on benchmarks incorporating algorithmic reasoning, assessing the reliability of these results necessitates an ability to cleanse models’ functional capabilities from memorization. In this paper, we propose an algorithmic benchmark comprising six tasks of infinite input domains where we can also disentangle and trace the correct, robust algorithm necessary for the task. This allows us to assess (i) models’ ability to extrapolate to unseen types of inputs, including new lengths, value ranges or input domains, but also (ii) to assess the robustness of the functional mechanism in recent models through the lens of their attention maps. We make the implementation of all our tasks and interoperability methods publicly available. [^1] https://github.com/michalspiegel/AttentionSpan
@@ 1 Introduction
The neural architecture of Transformer Vaswani et al. (2017) presents a backbone for a vast majority of modern language processing applications. A growing body of these applications, including code generation, conversational assistants, or data processing automatization, require Transformers to exhibit robust reasoning, i.e., an ability to identify and combine relevant pieces of information to infer new information Yu et al. (2024).
A critical component of the Transformer’s reasoning process is the Attention mechanism Bahdanau et al. (2014), which derives the representation of each newly generated token by weighing the representations of previous tokens.
@@
Despite the theoretical expressivity of Transformers in modelling even complex reasoning tasks Lin et al. (2021); Merrill and Sabharwal (2024), Transformers often depend on oversimplified, non-robust, or spurious features of data Mikula et al. (2024) causing even high-end models to fail in unexpected scenarios. This unreliability currently presents a critical bottleneck across a variety of applications. Bridging this gap requires fundamental improvements not only in architecture  Ye et al. (2025); Veličković et al. (2025) but also evaluation to rigorously assess how robust is the reasoning process of our models.
In this work, we contribute to bridging this gap by creating a new evaluation suite that presents a collection of diverse reasoning tasks.
Figure 1: Examples of reference attention maps we use to evaluate the models’ attention reasoning patterns. From top to bottom, you can see the attention scores with reference tokens highlighted in red, for addition, value assignment and FFLM task.
Each task has a solver algorithm, which generates a step-by-step solution and traces which past tokens are necessary for correctly generating the next one. This allows us to construct reference attention maps representing the ground truth reasoning patterns a successful model has to exhibit and compare it to the models’ actual attention map. We find that this comparison exposes some aspects of the model’s internal reasoning process, is highly predictive of the models’ reliability, and can identify sources of errors in the models’ reasoning process.
@@
We apply our benchmark to evaluate two different facets of generalization of existing language models: (i) an ability to learn to robustly execute the task in-distribution uncovering the robustness of models’ across perturbations of the input samples, and (ii) an ability to learn to generalize to out-of-distribution data, underlying the inherent limitations of existing architectures.
We find that Transformer models can learn to robustly execute algorithms on arbitrary in-distribution inputs. Based on our evaluation, models explain ID data by a seemingly correct algorithm, but despite that fail to generalize on new inputs.
When evaluating models on out-of-distribution data (OOD), especially longer inputs, models struggle to apply the reasoning attention pattern they learned in ID and make prediction errors rooted in not attending the correct tokens.
Our benchmark will empower future work in improving language models to not only assess the empirical improvements on our benchmarks but also to understand the implications of different architectural refinements on the robustness of models’ internal functioning cleansed from other covariates such as memorization.
@@ 2 Related Work
Closest to our work, CRLS-Text Markeeva et al. (2024) is a benchmark specialising in algorithmic reasoning implementing many traditional algorithms and trains and evaluates recent state-of-the-art LLMs. We build upon the methodology of CRLS-Text and extend it to allow for, not only accessing the performance, but also to provide means for interpretation and investigation of the results by means of the reference attention maps.
BIG-Bench Srivastava et al. (2023) is a massive benchmark comprised of more than 200 tasks, many of which specialize in evaluating algorithmic reasoning, e.g. addition or dyck languages. However, as a fixed test set, it is hard to use it to robustly evaluate models on extrapolation, while the recent work finds that BIG-Bench was indeed leaked into the training data of recent models Fajcik et al. (2024), including Qwen. We extend the tasks from BIG-Bench into configurable generators capable of generating infinite data, allowing training and evaluation while avoiding data contamination.
Flip-Flop Language Modeling is a synthetic task introduced by Liu et al. (2023). Authors introduce this simple algorithmic task to analyze hallucinations caused by attention glitches. We extend this idea and implement novel analysis of attention on a number of diverse algorithmic tasks.
@@ 3 AttentionSpan: Dataset and Evaluation Suite
To evaluate the reasoning robustness of Transformers, we introduce AttentionSpan, a framework for analyzing models’ attention patterns in step-by-step reasoning tasks.
AttentionSpan is composed of synthetic tasks with a highly controlled setting. Task instances (problems) can be randomly generated in arbitrary quantity and with configurable difficulty. The configuration also allows for systematic ID/IID splits that we also apply in our evaluations, including input lengths, ranges or domain. We detail provided configurations of AttentionSpan’s tasks in E.
Every problem has a single unambiguous solution, consisting of a deterministic sequence of steps that can be verified algorithmically.
A key contribution of our work is that every solution includes a reference attention mask that exactly specifies which past tokens are needed for correctly inferring the next one. It is important to mention that the reference maps are constructed in such way that they are independent of how the model implements the given algorithm. The indicated reference token are always crucial to completing the task. As we demonstrate in our experiments, reference attention masks are a powerful tool for inspecting the errors of transformers’ reasoning. We argue that they might facilitate future work in improving model reliability via architectural adjustments.
In the remainder of the Section, we describe the tasks in our suite. Examples of inputs and outputs can be found in Table 1.
@@
3.1 String Reversal
3.2 Long Multiplication
3.3 Long Addition
3.4 Successor
The Successor task requires a model to generate a sequence of natural numbers starting from a given initial value. It can be parametrized by the length of the series and the allowed range of the starting value. This is a straightforward task requiring precise representations of how digits form natural numbers.
3.5 Value Assignment
3.6 Flip Flop Language Modeling
@@ 4 Experiments and Evaluation
Model Type Task ID OOD
Acc. Attn Score Acc. Attn Score
From Scratch String Reversal 5.21 0.0578 0 0.0248
Long Addition 9.37 0.1713 0 0.0891
Long Multiplication 18 0.1302 0 0.0827
FFML 68.75 0.0129 50.2 0.0047
Value Assignment 4.17 0.3060 0 0.0683
Successor 100 0.4069 0 0.1450
Finetuned String Reversal 95.83 0.0836 53.83 0.0448
Long Addition 96.87 0.1380 1.61 0.0779
Long Multiplication 86 0.0432 0 0.0257
FFML 100 0.1854 99.2 0.1461
Value Assignment 100 0.1668 0 0.0378
Successor 100 0.4425 65.73 0.3770
Table 2: Performance of models trained from scratch and finetuned on various tasks. The changes in the mean Attn Score between ID and OOD are statistically significant in all cases.
Using the newly constructed benchmark, we aim to understand to what extent recent language models are capable of robustly representing and executing the underlying algorithms of our tasks. Towards this goal, we train and evaluate a popular LLama-3.2-1B-Instruct model on all tasks in two settings. First, we fine-tune the trained model with instruction in a few-shot setting. Secondly, we train the same architecture from scratch without instructions or few-shot examples. Differences in evaluations between the two variants can be attributed to the pre-training and instruction fine-tuning. Training setup and hyperparameter search are described in Appendix D.
We evaluate the models’ accuracy separately on ID and OOD data to measure their robustness to changes in problem length (See Appendix E). Next, we relate these results to the model’s ability to focus on relevant tokens, as provided by our dataset. To inspect which tokens the model considers in each reasoning step, we employ attention rollout Abnar and Zuidema (2020), a method for aggregating all models intermediate attention maps.
Using this aggregated attention map, we compute the proportion of attention scores allocated to tokens that our reference attention map identifies as necessary for correct prediction (See details in Appendix B). In particular, we measure and compare these metrics on the test instances where the model makes correct and erroneous predictions to uncover a possible pattern across error cases in the attention scores.
5 Results and Discusssion
Task OOD Acc. Attn Score (Correct) Attn Score (Error)
String Reversal
53.83
0.0455
0.0236
Successor
65.73
0.4141
0.3685
FFML
99.2
0.1948
0.3829
Table 3: Error in prediction significantly correlates with low attention score on reference tokens. The change in mean Attn Score is statistically significant in all cases.
Table 2 shows that models trained from scratch struggle with convergence and rarely generalize to OOD data, even with large sample sizes.
In contrast, an initialization from a pre-trained model dramatically improves the efficiency of convergence, achieving non-trivial performance already after using just tens or hundreds of samples. To a limited extent, resulting models can generalize to OOD data, even though the resulting accuracy consistently falls behind the ID performance.
We further analyze the distribution of attention scores on reference tokens. Welch’s t-test confirms a significant difference between ID and OOD data, with average attention scores dropping on OOD inputs. This may be due to longer sequences dispersing attention Veličković et al. (2025) or an inability to reliably identify key tokens.
Finally, in tasks where models perform well on in-distribution data, errors in OOD evaluations are often associated with a marked reduction in attention on reference tokens (see Table 3 and Appendix A). This pattern suggests one class of error, where insufficient attention directly contributes to faulty predictions. By contrast, tasks such as FFML show stable or even increased attention scores during errors, implying that different error mechanisms are at work. Importantly, because the OOD evaluation is not compromised by sequence length effects, the significant changes in attention scores for some tasks clearly reflect distinct classes of inference problems.
@@ 6 Conclusion
In this work, we introduced a novel algorithmic benchmark designed to assess both the extrapolation and reasoning capabilities of Transformer-based models robustly to memorization. Our evaluation framework, which leverages configurable tasks and reference attention maps, provides a transparent and fine-grained analysis of models’ internal reasoning processes beyond traditional accuracy metrics. This allows us to show that models trained from scratch struggle with generalization while pre-training projects into significant and consistent improvements on out-of-distribution inputs.
Importantly, our attention-based evaluation revealed that errors in reasoning are often associated with a dispersion of attention to relevant tokens. This finding not only validates our approach for diagnosing internal model behaviors but also lays the foundation for future architectural refinements aimed at enhancing robustness. By releasing our benchmark and methodologies, we hope to foster further research into the reliability and interpretability of Transformer models, ultimately contributing to the development of much-needed, robust AI systems for dynamic, real-world applications.
@@
Further, we acknowledge our focus only on a single model architecture as a limitation of our analyses. While at the time of writing, Llama model family represents a state-of-the-art among open-source models, we note that some trends, e.g. which tasks can/can not be learned can still be model-specific. Nevertheless, we focus our contribution instead to broadening a set of tasks and assuring a reliability of attention labels, leaving the investigation of further models to future work.
@@
Appendix A OOD Evaluation of String Reversal
Figure 2: In String Reversal, the model must learn a diagonal attention pattern. In ID evaluation (left), the model attributes high scores to all reference tokens. In OOD (right), it fails to do so for some tokens (highlighted in red), leading to prediction errors.

### v2

@@ Abstract
Can transformers learn to perform algorithmic tasks reliably across previously unseen input/output domains? While pre-trained language models show solid accuracy on benchmarks incorporating algorithmic reasoning, assessing the reliability of these results necessitates an ability to distinguish genuine algorithmic understanding from memorization. In this paper, we propose AttentionSpan, an algorithmic benchmark comprising five tasks of infinite input domains where we can disentangle and trace the correct, robust algorithm necessary for the task. This allows us to assess (i) models’ ability to extrapolate to unseen types of inputs, including new lengths, value ranges or input domains, but also (ii) to assess the robustness of their learned mechanisms. By analyzing attention maps and performing targeted interventions, we show that attention mechanism directly causes failures in extrapolation. We make the implementation of all our tasks and interpretability methods publicly available. [^1] https://github.com/michalspiegel/AttentionSpan
@@ 1 Introduction
The neural architecture of Transformer Vaswani et al. (2017) presents a backbone for a vast majority of modern language processing applications. A growing body of these applications, including code generation, conversational assistants, or data processing automation, requires Transformers to exhibit robust reasoning, i.e., an ability to identify and combine relevant pieces of information to infer new information Yu et al. (2024).
@@
Transformers rely on the Attention mechanism Bahdanau et al. (2014) as their only means of mixing information across token streams, which is essential for long-form reasoning. However, its effectiveness often degrades, especially with increasing sequence lengths Veličković et al. (2025). Despite the strong theoretical expressivity Yun et al. (2020); Merrill and Sabharwal (2024), Transformers often depend on spurious features of data Mikula et al. (2024), causing even high-end models to fail in unexpected scenarios. This unreliability currently presents a critical bottleneck across a variety of applications. Bridging this gap requires fundamental improvements not only in architecture  Ye et al. (2025); Veličković et al. (2025) but also evaluation to rigorously assess how robust the reasoning process of our models is.
In this work, we contribute to bridging this gap by creating AttentionSpan, a new evaluation suite focused on assessing fundamental reasoning capabilities of language models in out-of-distribution scenarios.
Figure 1: Examples of attention patterns that we use to evaluate the robustness of models’ reasoning. Each grid displays normalized attention scores (post-softmax), aggregated across heads and layers via attention rollout, overlaid with a reference map (in red) denoting necessary input tokens. Visualizations show the tasks of addition, value assignment, and FFML (§3; ordered top to bottom, respectively).
Each task in AttentionSpan has a solver algorithm, which generates a step-by-step solution, together with a trace of which past tokens are necessary for correctly generating the next one. This enables deeper analyses of the model’s reasoning behavior, not possible with previous resources:
  • Provision of reference attention masks representing the ground-truth reasoning patterns that a successful model has to follow in order to achieve a correct prediction.
  • Control over distribution shifts, i.e., systematic changes in the constructed dataset that allow for a reliable assessment of models’ reasoning in novel scenarios.
@@
First, we apply our benchmark to evaluate two different facets of generalization of existing language models: (i) an ability to learn to accurately combine the information necessary for the task in-distribution, uncovering the models ability to fit an algorithm, and (ii) an ability to generalize to out-of-distribution data, exploiting the inherent limitations of existing architectures; We find that Transformer models can learn to explain in-distribution (ID) data by a seemingly correct algorithm, but despite that, fail to generalize to unseen, out-of-distribution (OOD) inputs.
Finally, we hypothesize that out-of-distribution (OOD) errors stem from attention misalignments and verify this by an intervention experiment that directly manipulates attention to match reference attention maps from AttentionSpan. We find that such intervention increases models’ OOD accuracy by up to 90 percentage points, and provides consistent improvements even with increasing input sequence lengths. This establishes a direct causal link between attention failures and poor generalization, pinpointing attention as a key bottleneck for length extrapolation.
The presented benchmark empowers future work by a toolset for further, deeper analyses of models’ internal functioning, isolated from other confounders such as memorization. Our findings also motivate future work in architectural refinements, particularly those addressing limitations of the current Attention mechanism.
@@ 2 Related Work
Closely related to our work, CRLS-Text Markeeva et al. (2024) is a benchmark for algorithmic reasoning that implements traditional algorithms to train and evaluate state-of-the-art LLMs. We build upon the methodology of CRLS-Text and extend it to allow for, not only accessing the performance, but also to provide means for interpretation and investigation of the results by means of the reference attention masks.
BIG-Bench Srivastava et al. (2023) is a massive benchmark comprised of more than 200 tasks, many of which specialize in evaluating algorithmic reasoning, e.g. addition or dyck languages. However, as a fixed test set, it is hard to use it to robustly evaluate models on extrapolation, while the recent work finds that BIG-Bench was indeed leaked into the training data of recent models Fajcik et al. (2024), including Qwen. We extend the tasks from BIG-Bench into configurable generators capable of generating infinite data, enabling a controlled training and evaluation setting, avoiding data contamination.
Flip-Flop Language Modeling is a synthetic task introduced by Liu et al. (2023). The authors introduce this simple algorithmic task to analyze hallucinations caused by attention glitches. We extend this idea and implement novel analysis of attention on a number of diverse algorithmic tasks.
@@ 3 AttentionSpan: Dataset and Evaluation Suite
To evaluate the reasoning robustness of Transformers, we introduce AttentionSpan, a benchmark for analyzing models’ attention patterns in step-by-step reasoning tasks.
AttentionSpan is composed of synthetic tasks allowing for a fully controlled generation setting and guaranteed out-of-distribution assessment. The tasks include string reversal, addition, multiplication, flip-flop language modeling Liu et al. (2023), and value assignment, all described in detail in Appendix A. Examples of inputs and outputs can be found in Table 1. Task instances (problems) can be randomly generated in arbitrary quantities and with configurable difficulty. The configuration also allows for systematic ID-OOD splits applied in our evaluations, including input lengths, ranges or domains. We detail our evaluated ID-OOD configurations of AttentionSpan’s tasks in Appendix E.
@@
3.1 Reference Attention Masks
A key novelty of AttentionSpan is the provision of a reference attention mask with every generated data sample. This mask precisely identifies the past tokens essential for correct next-token inference. We utilize these masks as an expected attention pattern for an ideal model and subsequently measure the alignment of a model’s learned attention with this reference. Our experiments demonstrate that reference attention masks are a powerful tool for diagnosing reasoning errors in Transformers, evidencing that they could facilitate future research aimed at enhancing model reliability, e.g., through architectural modifications.
The reference attention mask is a discrete boolean matrix. For each target token to be predicted, it identifies critical (reference) past tokens that the model should attend to for robust problem-solving (see tokens highlighted in red in Figure 1). An element in the matrix is set to 1 if the corresponding past token carries information relevant to predicting the target token; otherwise, it is set to 0, indicating irrelevant tokens that may not be attended for the current prediction.
Model Task ID Acc. OOD Acc. OOD Partial Acc.
Llama-3.2 String Reversal 95.83 53.83 96.18
Long Addition 96.87 1.61 64.76
Long Multiplication 86.00 0.00 73.87
FFML 100.00 99.20 99.79
Value Assignment 93.95 0.52 65.26
Qwen2.5 String Reversal 98.95 21.77 76.26
Long Addition 100.00 44.75 92.77
Long Multiplication 56.25 0.00 80.14
FFML 100.00 88.50 96.71
Value Assignment 76.04 1.04 81.06
gemma-3 String Reversal 96.87 6.04 37.64
Long Addition 100.00 2.62 67.85
Long Multiplication 89.58 0.00 77.16
FFML 100.00 90.12 96.71
Value Assignment 98.95 0.00 19.35
Table 2: Accuracy of finetuned models on AttentionSpan tasks with consistent in-distribution and out-of-distribution splits. Despite a sharp decline in OOD Accuracy in almost all cases, the OOD Partial Accuracy reveals that models correctly predict a large proportion of target tokens, indicating some extrapolation abilities are present.
Model          Task Mean AttnScore (Correct) (%) Mean AttnScore (Error) (%)
Llama-3.2 String Reversal $4.55\pm 0.02$ $2.36\pm 0.09$
Value Assignment $3.33\pm 0.03$ $1.26\pm 0.02$
Qwen2.5 String Reversal $3.07\pm 0.06$ $2.17\pm 0.05$
Value Assignment $1.27\pm 0.03$ $1.20\pm 0.06$
gemma-3 String Reversal $3.97\pm 0.12$ $2.30\pm 0.03$
Value Assignment $5.04\pm 0.12$ $4.91\pm 0.04$
Table 3: Errors in prediction are associated with lower attention score. We find a statistically significant difference (Welch’s t-test) between attention scores on correct and incorrect (target token) predictions.
@@ 4 Experiments and Evaluation
Figure 2: We demonstrate that by intervening on models’ (Llama-3.2-1B-Instruct, Qwen2.5-1.5B-Instruct) activations and directly adjusting their attention scores to reinforce our reference attention pattern, we are able to drastically improve the length extrapolation performance over the vanilla models without invervention, attributing the failure to extrapolate to the attention mechanism.
Building on AttentionSpan, we assess how robustly recent language models can learn and execute the underlying algorithm. Towards this goal, we explore models in two distinct settings: (1) training models from random initialization, and (2) fine-tuning pre-trained models – specifically Llama-3.2-1B-Instruct Grattafiori et al. (2024), Qwen2.5-1.5B-Instruct Yang et al. (2024), and gemma-3-1b-it Team et al. (2025) – on all tasks using a few-shot instruction prompt, in order to evaluate the contribution of pre-training. To evaluate these models, we use AttentionSpan to introduce systematic distribution shifts, altering the length of input sequences and other key parameters (for details on ID/OOD settings see Appendix E). Training setup and hyperparameter search are further detailed in Appendix D.
In addition to strict exact-match accuracy of whole output sequences, we measure Partial Accuracy as the fraction of correct next-token predictions in the target sequence:
Partial $\text{Partial Acc}=\frac{1}{N}\sum_{i=1}^{N}\frac{1}{|T_{i}|}\sum_{j=1}^{|T_{i}|}\mathbb{I}(\hat{y}_{i,j}=y_{i,j})$ (1)
where $N$ is the total number of samples, $T_{i}$ is the target sequence for sample $i$, $\hat{y}_{i,j}$ and $y_{i,j}$ are the predicted and target tokens at position $j$ respectively, and $\mathbb{I}(\cdot)$ is the indicator function.
We use the reference attention masks to track which tokens the model considers at each reasoning step. We rely on attention rollout Abnar and Zuidema (2020), a standard method for aggregating attention across heads and layers. This allows us to visualize the attention patterns (e.g., Figure 1) and assess whether the model learned the expected attention patterns in a generalized fashion. The formula for attention rollout (as defined in Abnar and Zuidema (2020)) can be expressed as a recursive product:
$R=\prod_{l=1}^{L}\left(\frac{1}{H}\sum_{h=1}^{H}A_{l,h}+I\right)$ (2)
where $R$ is the final attention rollout matrix, $L$ is the total number of self-attention layers, $H$ is the number of attention heads in each layer, $A_{l,h}$ is the (post-softmax) attention score matrix for head $h$ in layer $l$, $I$ is the identity matrix, which accounts for the residual connections.
Quantitatively, we leverage our dataset’s reference attention masks and aggregated attention scores to calculate the proportion of attention scores assigned to tokens identified as essential for correct prediction, which we refer to as Attn Score in Table 3 (see Appendix B for details). Subsequently, through the lens of Attn Score, we explore differences in models’ behavior on correct and incorrect (token-level) predictions to identify systematic attention patterns associated with errors. Formally, Attn Score is defined as:
Attn $\text{Attn Score}\!=\!\frac{1}{N}\sum_{i=1}^{N}\!\!\left(\!\frac{1}{|T_{i}|}\sum_{j=1}^{|T_{i}|}\!\left(\sum_{k\in S_{i}}a_{i,j,k}\!\right)\!\!\right)$ (3)
where $N$ is the total number of samples in the dataset, $T_{i}$ is the sequence of target (output) tokens for sample $i$, and $|T_{i}|$ is its length, $S_{i}$ is the set of indices corresponding to the reference input tokens for sample $i$, $a_{i,j,k}$ is the normalized attention score in aggregated $R$ (Eq. (2)) from the $j$-th output token to the $k$-th input token in sample $i$. We will represent AttnScore as percentages.
5 Results: Attention as the Bottleneck
As detailed in Table 4, the fine-tuned models exhibit difficulty generalizing to OOD data (See Appendix E for details on exact ID/OOD parameters). However, their ability to correctly predict a majority of target tokens in each sample (demonstrated by high partial accuracy) suggests underlying length extrapolation capabilities and significant potential for future improvement.
Models trained from random initialization, despite fitting the in-distribution (ID) data, exhibited near-zero accuracy in OOD evaluations (see Appendix G). Given this lack of generalization, we focused our subsequent analysis on fine-tuned models, with the full results from this initial experiment available in Appendix G. This outcome suggests that the pre-training phase is advantageous for generalization and convergence.
Our analyses reveal that for String Reversal and Value Assignment, out-of-distribution (OOD) errors are linked to reduced attention on reference tokens (Table 3). The appearance of this phenomenon in these specific tasks is significant, as they are representative of copying tasks—a recently highlighted benchmark class Arjovsky et al. (2016); Jelassi et al. (2024)—and embody two distinct, fundamental retrieval modes: positional (e.g., sorting, reordering) and content-based (e.g., using variables or a knowledge base). This pattern suggests that insufficient attention to reference tokens contributes to faulty predictions. We observe this issue across diverse pre-trained models and architectures, indicating a naturally emergent, general problem.
5.1 Attention Misallocation as a Failure Mode
To better understand the nature of these OOD failures, we visualize the attention patterns of individual heads that specialize in the required lookup algorithm. Figure 3 demonstrates how attention scores scatter away from the reference diagonal as input sequence length increases. Table 3 quantifies this, showing that prediction errors occur precisely when attention deviates from the reference diagonal. Interestingly, we observe that attention does not simply diffuse across neighboring tokens in OOD scenarios. Instead, it remains sharp but incorrectly shifts to a distant, irrelevant token. This suggests a potential failure in how positional embeddings like RoPE generalize to long-range positional relationships not seen during training.
Figure 3: Attention from a single head for String Reversal identified as performing our reference diagonal lookup algorithm. As sequence length increases from in-distribution (top) to out-of-distribution (bottom), attention scatters from the correct diagonal. This misallocation (highlighted in red) directly causes prediction errors.
5.2 Attention Reinforcement
To causally validate that insufficient attention to reference tokens drives OOD failures, we first identified key attention heads implementing our expected reference attention pattern. To do this, we select the top 10% of the heads with the largest sum of attention scores on reference tokens in ID setting. We then selectively reinforce the reference attention in OOD inference by directly increasing the post-softmax attention to reference tokens of such-identified heads. As shown in Figure 2, this intervention leads to an absolute increase in accuracy on OOD samples of up to 90 percentage points, consistently across all input lengths. This provides causal evidence that insufficient attention to reference tokens largely contributes to extrapolation failures. For details on the implementation of the intervention, see Appendix F.
While our intervention targets attention outputs, we hypothesize the root cause is the extrapolation failure of positional embeddings like RoPE. The key-query similarity function is learned over a limited range of relative distances. For longer OOD sequences, the model encounters distances outside this familiar range, causing the learned function to fail and produce the unstable, misallocated attention scores we observe.
@@ 6 Conclusion
We introduced AttentionSpan, a novel algorithmic reasoning benchmark including reference attention masks, allowing assess Transformer extrapolation under different, parametrized distribution shifts, thus minimizing memorization effects. Our evaluations reveal that while pre-training helps, models struggle with out-of-distribution (OOD) generalization even on simple algorithmic tasks. Through attention analysis and targeted interventions, we causally attribute these OOD failures to Attention’s inability to robustly identify essential tokens. We show that reinforcing correct attention patterns significantly improved OOD accuracy and length extrapolation, pinpointing attention as a critical bottleneck.
@@
Appendix A Task Descriptions
A.1 String Reversal
A.2 Long Multiplication
A.3 Long Addition
A.4 Value Assignment
A.5 Flip Flop Language Modeling